# Entropy Meets Importance: A Unified Head Importance–Entropy Score for Stable and Efficient Transformer Pruning

### v1

@@ 5.3.1 Head Removal Patterns (Heatmap)
Pruning heatmap analysis provides empirical support for this distinction. As shown in Fig. 3, HIS (left panels) tends to remove heads primarily from the lower layers, producing an approximately bottom-up pattern. This behavior is guided by a one-step gradient saliency, which estimates the importance of each attention head based on a single backward pass through the model— this assigns higher importance to heads whose activations have a larger immediate effect on the loss. Meanwhile, HIES (right panels) yields a more dispersed selection spanning lower, middle, and upper layers. We attribute this to the entropy-aware term, which leverages structural properties of the attention distribution (concentration vs. dispersion) in addition to gradient sensitivity, thereby promoting diversity across layers in pruning decisions. Consequently, HIES exhibits more stable performance across pruning ratios, yielding flatter accuracy–sparsity curves than HIS. Additional results are provided in Appendix D.2.
@@ 5.3.3 Generalizability to Attention Variants and Tasks
First, across different attention-based variants, including ViT and LLaVA, HIES consistently shifts the region of sharp accuracy drop to more aggressive sparsity levels compared to HIS. In particular, at high pruning ratios where HIS accuracy collapses, HIES achieves 73.17% higher accuracy on $\text{ViT}_{\text{Large}}$ at 30% sparsity, compared to HIS (with 12.80% of accuracy). This indicates that HIES effectively captures structural importance regardless of the specific model configuration, enabling stable and reliable pruning.
Second, the advantages of HIES are further extended on complex multi-modal tasks. HIES shows improvement of 11.71% at 50% sparsity compared to HIS on VizWiz-VQA and MM-Vet. HIS frequently suffers from sharp accuracy drops on these tasks, as it relies solely on gradient-based head importance and fails to account for the structural diversity of attention heads. As a result, pruning based on HIS alone may remove heads that are essential for preserving cross-modal alignment and semantic grounding. In contrast, HIES leverages AE to capture these structural signals, resulting in more stable performance across challenging vision-language benchmarks. HIES also demonstrates robust and stable performance on downstream tasks (i.e., CIFAR-100, Food-101, and Fashion MNIST) — detailed results are provided in Appendix D.4.
@@ Appendix A Related Work
Model Compression. Language models [Devlin et al., 2018, Liu et al., 2019, Lewis et al., 2019] have rapidly advanced in scale and capability, intensifying the demand to reduce their parameter sizes and inference latency [Molchanov et al., 2017, Gordon et al., 2020]. To compress large language models efficiently, various approaches have been explored, including pruning [Liu et al., 2021b, Kurtic et al., 2022, Xu et al., 2021], knowledge distillation [Sun et al., 2019, 2020, Pan et al., 2020], quantization [Bai et al., 2020, Yao et al., 2022, Zafrir et al., 2019, Chen et al., 2025a], and other techniques, like low-rank approximation methods [Saha et al., 2024, Li et al., 2023, Wong et al., 2025] or weight space decomposition methods such as [Ashkboos et al., 2024].
Structured Pruning. Among these, structured pruning—which removes entire architectural components rather than individual weights— can be performed at various granularities, such as whole layers [Fan et al., 2020], multi-head attention modules [Michel et al., 2019, Voita et al., 2019], or feed-forward networks [Lagunas et al., 2021, Hou et al., 2020b]. Recently, attention head pruning has gained particular traction in the context of large language models. It directly reduces attention FLOPs and KV-cache size by lowering the number of active heads while preserving the layer topology, thereby simplifying checkpoint compatibility and serving integration. Consequently, several large-scale studies adopt head-level pruning as a practical axis in LLM compression pipelines [Ma et al., 2023, Jaradat and et al., 2024, Muralidharan et al., 2024]. Ma et al. [2023] propose a unified framework that integrates structured head pruning into the training pipeline of large language models, achieving substantial sparsity without accuracy degradation.
Entropy in Reinforcement Learning and Transformers. Entropy has long been employed in reinforcement learning (RL) as a means of encouraging exploration and preserving policy diversity. By introducing an entropy term into the policy objective, RL methods prevent premature convergence to deterministic strategies and mitigate policy collapse [Bharadhwaj et al., 2020, Liu et al., 2021a, Wang et al., 2025]. Extensions of this idea include parameter-space entropy regularization to explicitly control diversity [Han et al., 2023], large-deviation interpretations of entropy-regularized RL [Arriojas et al., 2023], and state-distribution entropy regularization for improved robustness and generalization [Ashlag et al., 2025].
@@ Appendix B Proofs and Details
Overview. This appendix collects the formal analyses that underpin Section 4.2 and fixes notation and technical conventions used throughout. We first develop loss-increase control for head pruning via gradient-based head importance (HIS), derive operator-norm curvature bounds for the cross-entropy objective, and justify the first-order approximation by quantifying when quadratic terms are negligible (Appendix B.1).
@@ B.1.6 Remarks on HIS with Absolute Values
The absolute value in equation 1 is part of the definition to prevent cancellation across samples; consequently, the triangle inequality turns the first-order term into an additive upper bound $\sum_{h}(1-m_{h})\,\mathrm{HIS}_{h}$ (cf. Appendix B.1.1). If $\langle\nabla_{A_{h}}\mathcal{L},A_{h}\rangle_{F}<0$ on some samples, masking that head could locally decrease the loss; the metric remains conservative by construction.
@@ B.2.4 Entropy–Total Variation (TV) Control
Lemma 4 (Entropy–TV inequality).
@@
Applying Lemma 4 to equation 9 token-wise and averaging,
@@ B.2.5 Stability and Generalization
Let $\gamma:=L_{\ell}\,\mathbb{E}_{\mathcal{S}}[\|\Delta(x)\|_{2}]$. By on-average replace-one stability [Bousquet and Elisseeff, 2002, Def. 6],
@@
Using equation 10 and Jensen for $\sqrt{\cdot}$,
@@
Taking a representative $n$ (e.g., average/max effective length) yields the main-text constant $C_{\mathrm{AE}}=\sqrt{8}\,M\,\sqrt{|\text{H}|\rho\,\log n}$ and Eq. equation 4.
@@ B.2.6 Constants and Practical Remarks
Let $\operatorname{supp}(m):=\{h:\,m_{h}=1\}$ denote the set of retained heads. Suppose an admissible mask $m^{\prime}$ with $|\operatorname{supp}(m^{\prime})|=k$ is not optimal. Then there exist $i\in\operatorname{supp}(m^{\prime})$ and $j\notin\operatorname{supp}(m^{\prime})$ such that $\text{HIES}_{j}>\text{HIES}_{i}$. Consider the mask $\tilde{m}$ that swaps $i$ and $j$ (retain $j$, prune $i$); the constraint in equation 6 is preserved. The objective in equation 5 changes by
@@ C.4.1 Natural Language Understanding Task
Table 5: Summary of the NLG benchmark.
@@ C.4.2 Image Classification Task
C.4.3 Multi-modal Vision-Language Task
To evaluate the effectiveness of our pruning method on multi-modal vision-language models (VLMs), we used two benchmark datasets: VizWiz-VQA [Gurari et al., 2018] and MM-Vet [Yu et al., 2024]. The evaluation was conducted using the $\text{LLaVA1.5}_{\text{7B}}$ model.
@@ D.1 Orthogonality Analysis
We provide an empirical sanity check supporting the assumptions above. Experiments use TinyBERT on SST-2 (Vishnou/TinyBERT_SST2). For each head we compute layerwise-normalized HIS and attention-entropy (AE) scores, stack them into vectors $u$ and $v$, and form the centered versions $\tilde{u}$ and $\tilde{v}$ by subtracting each vector’s mean (a finite-sample proxy for projection onto the zero-sum subspace). The following sample statistics were obtained:
@@
The covariance magnitude is small on this batch, indicating weak coupling between the two directions and lending empirical support to the “(near) uncorrelatedness” assumption. We recommend reporting the same diagnostics averaged over multiple batches to reduce sampling noise and to provide confidence intervals.
D.2 Heatmap of Importance Scores and Pruning Results

### v2

@@ 5.3.1 Head Removal Patterns (Heatmap)
Pruning heatmap analysis provides empirical support for this distinction. As shown in Figure 3, HIS (left panels) tends to remove heads primarily from the lower layers, producing an approximately bottom-up pattern. This behavior is guided by a one-step gradient saliency, which estimates the importance of each attention head based on a single backward pass through the model— this assigns higher importance to heads whose activations have a larger immediate effect on the loss. Meanwhile, HIES (right panels) yields a more dispersed selection spanning lower, middle, and upper layers — other tasks also exhibit similar patterns (details are provided in Appendix D.4). We attribute this to the entropy-aware term, which leverages structural properties of the attention distribution (concentration vs. dispersion) in addition to gradient sensitivity, thereby promoting diversity across layers in pruning decisions. Consequently, HIES exhibits more stable performance across pruning ratios, yielding flatter accuracy–sparsity curves than HIS.
@@ 5.3.3 Generalizability to Attention Variants and Tasks
First, across different attention-based variants, including ViT and LLaVA, HIES consistently shifts the region of sharp accuracy drop to more aggressive sparsity levels compared to HIS. In particular, at high pruning ratios where HIS accuracy collapses, HIES achieves 49.82% higher accuracy on $\text{ViT}_{\text{Large}}$ at 20% sparsity, compared to HIS (with 55.20% of accuracy). This indicates that HIES effectively captures structural importance regardless of the specific model configuration, enabling stable and reliable pruning.
Second, the advantages of HIES are further extended on complex multi-modal tasks. HIES shows improvement of 11.71% at 50% sparsity compared to HIS on VizWiz-VQA and MM-Vet. HIS frequently suffers from sharp accuracy drops on these tasks, as it relies solely on gradient-based head importance and fails to account for the structural diversity of attention heads. As a result, pruning based on HIS alone may remove heads that are essential for preserving cross-modal alignment and semantic grounding. In contrast, HIES leverages AE to capture these structural signals, resulting in more stable performance across challenging vision-language benchmarks.
Third, we evaluate HIES on large-scale reasoning benchmarks, including GSM8K and MMLU, to assess its robustness beyond classification-style tasks. The results show consistent gains under reasoning-intensive settings and are reported in Appendix D.6. We also report downstream evaluations on CIFAR-100, Food-101, and Fashion-MNIST, demonstrating that the benefits of HIES extend to diverse vision benchmarks. Detailed results are provided in Appendix D.7.
@@ 6 Conclusion
Impact Statements
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix
A. Related Work p. A
B. Proofs and Details p. B
B.1. Loss-increase Control via Head Importance (Section 4.2.1) ........................................................................................................................................................ p. B.1
B.1.1. Loss Bound with Operator-Norm Control ........................................................................................................................................................ p. B.1.1
B.1.2. Norms and Inner Products ........................................................................................................................................................ p. B.1.2
B.1.3. Estimating $\|W^{O}\|_{2}$ and Blockwise Norms via Power Iteration ........................................................................................................................................................ p. B.1.3
B.1.4. Cross-Entropy Curvature and Propagation to $\mathbf{y}$ ........................................................................................................................................................ p. B.1.4
B.1.5. Why the Quadratic Term Is Typically Negligible ........................................................................................................................................................ p. B.1.5
B.1.6. Remarks on HIS with Absolute Values ........................................................................................................................................................ p. B.1.6
B.2. Generalization Gap and Attention Entropy (Section 4.2.2) ........................................................................................................................................................ p. B.2
B.2.1. Notation and Token Averaging ........................................................................................................................................................ p. B.2.1
B.2.2. Neighboring Datasets and Why They Appear ........................................................................................................................................................ p. B.2.2
B.2.3. From Attention Perturbation to Output Perturbation ........................................................................................................................................................ p. B.2.3
B.2.4. Entropy–Total Variation (TV) Control ........................................................................................................................................................ p. B.2.4
B.2.5. Stability and Generalization ........................................................................................................................................................ p. B.2.5
B.2.6. Constants and Practical Remarks ........................................................................................................................................................ p. B.2.6
B.3. Risk Upper Bound and HIES Minimization (Section 4.2.3) ........................................................................................................................................................ p. B.3
B.4. Orthogonality and Complementarity (Section 4.2.4) ........................................................................................................................................................ p. B.4
C. Experimental Setup p. C
C.1. Experimental Setup for Motivation Study ........................................................................................................................................................ p. C.1
C.2. Model ........................................................................................................................................................ p. C.2
C.3. Computing Resources ........................................................................................................................................................ p. C.3
C.4. Dataset Statistics ........................................................................................................................................................ p. C.4
C.4.1. Natural Language Understanding Task ........................................................................................................................................................ p. C.4.1
C.4.2. Image Classification Task ........................................................................................................................................................ p. C.4.2
C.4.3. Reasoning Task ........................................................................................................................................................ p. C.4.3
C.4.4. Multi-modal Vision-Language Task ........................................................................................................................................................ p. C.4.4
D. Additional Experimental Results p. D
D.1. Orthogonality Analysis ........................................................................................................................................................ p. D.1
D.2. Seed Sensitivity and Statistical Significance ........................................................................................................................................................ p. D.2
D.3. Effect of Calibration Dataset Size ........................................................................................................................................................ p. D.3
D.4. Heatmap of Importance Scores and Pruning Results ........................................................................................................................................................ p. D.4
D.5. 3D Analysis of Attention Head Importance Scores ........................................................................................................................................................ p. D.5
D.5.1. Difference in Pruning Patterns ........................................................................................................................................................ p. D.5.1
D.5.2. Performance Inversion Across Sparsity Regimes ........................................................................................................................................................ p. D.5.2
D.6. Experimental Results on Large-scale Reasoning Tasks ........................................................................................................................................................ p. D.6
D.7. Experimental Results on Downstream Tasks ........................................................................................................................................................ p. D.7
D.8. Sensitivity Analysis - Ablation on $\alpha$ ........................................................................................................................................................ p. D.8
D.9. Combining Attention Entropy with Other Importance Signals ........................................................................................................................................................ p. D.9
E. Efficiency Analysis p. E
E.1. Computational Efficiency and FLOPs Reduction ........................................................................................................................................................ p. E.1
E.2. Runtime Overhead and Implementation Details ........................................................................................................................................................ p. E.2
@@ Appendix A Related Work
Model Compression. Language models [Devlin et al., 2019, Liu et al., 2019, Lewis et al., 2020] have rapidly advanced in scale and capability, intensifying the demand to reduce their parameter sizes and inference latency [Molchanov et al., 2017, Gordon et al., 2020]. To compress large language models efficiently, various approaches have been explored, including pruning [Liu et al., 2021, Kurtic et al., 2022, Xu et al., 2021], knowledge distillation [Sun et al., 2019, 2020, Pan et al., 2021], quantization [Bai et al., 2021, Yao et al., 2022, Zafrir et al., 2019, Chen et al., 2025a], and other techniques, like low-rank approximation methods [Saha et al., 2024, Li et al., 2023b, Wong et al., 2025] or weight space decomposition methods such as [Ashkboos et al., 2024].
Structured Pruning. Among these, structured pruning—which removes entire architectural components rather than individual weights— can be performed at various granularities, such as whole layers [Fan et al., 2019], multi-head attention modules [Michel et al., 2019, Voita et al., 2019], or feed-forward networks [Lagunas et al., 2021, Hou et al., 2020]. Recently, attention head pruning has gained particular traction in the context of large language models. It directly reduces attention FLOPs and KV-cache size by lowering the number of active heads while preserving the layer topology, thereby simplifying checkpoint compatibility and serving integration. Consequently, several large-scale studies adopt head-level pruning as a practical axis in LLM compression pipelines [Ma et al., 2023, Jaradat et al., 2024, Muralidharan et al., 2024]. Ma et al. [2023] propose a unified framework that integrates structured head pruning into the training pipeline of large language models, achieving substantial sparsity without accuracy degradation.
Entropy in Reinforcement Learning and Transformers. Entropy has long been employed in reinforcement learning (RL) as a means of encouraging exploration and preserving policy diversity. By introducing an entropy term into the policy objective, RL methods prevent premature convergence to deterministic strategies and mitigate policy collapse [Bharadhwaj et al., 2020, Xiao et al., 2021, Wang et al., 2025]. Extensions of this idea include parameter-space entropy regularization to explicitly control diversity [Han et al., 2023], large-deviation interpretations of entropy-regularized RL [Arriojas et al., 2023], and state-distribution entropy regularization for improved robustness and generalization [Ashlag et al., 2025].
@@ Appendix B Proofs and Details
Overview. This appendix collects the formal analyses that underpin Section 4.2 and fixes notation and technical conventions used throughout. We first develop loss-increase control for head pruning via gradient-based HIS, derive operator-norm curvature bounds for the cross-entropy objective, and justify the first-order approximation by quantifying when quadratic terms are negligible (Appendix B.1).
@@ B.1.6 Remarks on HIS with Absolute Values
The absolute value in (1) is part of the definition to prevent cancellation across samples; consequently, the triangle inequality turns the first-order term into an additive upper bound $\sum_{h}(1-m_{h})\,\mathrm{HIS}_{h}$ (cf. Appendix B.1.1). If $\langle\nabla_{A_{h}}\mathcal{L},A_{h}\rangle_{F}<0$ on some samples, masking that head could locally decrease the loss; the metric remains conservative by construction.
@@ B.2.4 Entropy–Total Variation (TV) Control
Lemma B.1 (Entropy–TV inequality).
@@
Applying Lemma B.1 to (10) token-wise and averaging,
@@ B.2.5 Stability and Generalization
Let $\gamma:=L_{\ell}\,\mathbb{E}_{\mathcal{S}}[\|\Delta(x)\|_{2}]$. By on-average replace-one stability [Bousquet and Elisseeff, 2002],
@@
Using (11) and Jensen for $\sqrt{\cdot}$,
@@
Taking a representative $n$ (e.g., average/max effective length) yields the main-text constant $C_{\mathrm{AE}}=\sqrt{8}\,M\,\sqrt{|\text{H}|\rho\,\log n}$ and Eq. (5).
@@ B.2.6 Constants and Practical Remarks
Let $\operatorname{supp}(m):=\{h:\,m_{h}=1\}$ denote the set of retained heads. Suppose an admissible mask $m^{\prime}$ with $|\operatorname{supp}(m^{\prime})|=k$ is not optimal. Then there exist $i\in\operatorname{supp}(m^{\prime})$ and $j\notin\operatorname{supp}(m^{\prime})$ such that $\text{HIES}_{j}>\text{HIES}_{i}$. Consider the mask $\tilde{m}$ that swaps $i$ and $j$ (retain $j$, prune $i$); the constraint in (7) is preserved. The objective in (6) changes by
@@ C.4.1 Natural Language Understanding Task
Table 5: Summary of the NLU benchmark.
@@ C.4.2 Image Classification Task
C.4.3 Reasoning Task
To evaluate the effectiveness of our pruning method on reasoning tasks, we use two benchmark datasets: GSM8K [Cobbe et al., 2021] and MMLU [Hendrycks et al., 2020].
GSM8K: is a dataset of 8.5K high quality, linguistically diverse grade school math word problems. GSM8K supports the task of question answering on basic mathematical problems requiring multi-step reasoning.
MMLU: is a massive multitask benchmark consisting of multiple-choice questions spanning diverse domains, including the humanities, social sciences, and hard sciences. The benchmark covers 57 tasks such as elementary mathematics, U.S. history, computer science, and law, and requires models to possess broad world knowledge and strong problem-solving abilities to achieve high performance.
By using both GSM8K and MMLU, we evaluate our pruning methods from complementary perspectives: GSM8K assesses mathematical reasoning under multi-step computation, while MMLU measures the preservation of broad world knowledge and multi-domain problem-solving ability.
C.4.4 Multi-modal Vision-Language Task
To evaluate the effectiveness of our pruning method on multi-modal vision-language models (VLMs), we use two benchmark datasets: VizWiz-VQA [Gurari et al., 2018] and MM-Vet [Yu et al., 2023]. The evaluation was conducted using the $\text{LLaVA1.5}_{\text{7B}}$ model.
@@ D.1 Orthogonality Analysis
We provide an empirical sanity check supporting the Lemma 4.3. Experiments use TinyBERT on SST-2 (Vishnou/TinyBERT_SST2). For each head we compute layerwise-normalized HIS and AE scores, stack them into vectors $u$ and $v$, and form the centered versions $\tilde{u}$ and $\tilde{v}$ by subtracting each vector’s mean (a finite-sample proxy for projection onto the zero-sum subspace). The following sample statistics were obtained:
@@
The covariance magnitude is small on this batch, indicating weak coupling between the two directions and lending empirical support to the “(near) uncorrelatedness” assumption.
We further extend the orthogonality analysis across models, with results reported in Table 7 and Table 8.
Table 7: Orthogonality analysis using $\text{BERT}_{\text{base}}$ on the GLUE benchmark.
Task $\tilde{\text{u}}] E$$\tilde{\text{v}}$] Cov($\tilde{\text{u}}$$\tilde{\text{v}}$) E[$\tilde{\text{u}}$$\tilde{\text{v}}$⟩] Correlation
COLA 6.94E-03 4.08E-01 -1.40E-05 2.82E-03 -1.54E-02
SST2 6.94E-03 4.42E-01 -3.60E-05 3.03E-03 -3.41E-02
MRPC 6.94E-03 5.27E-01 2.41E-04 3.90E-03 2.32E-01
STSB 6.94E-03 4.30E-01 -1.96E-04 2.79E-03 -1.65E-01
QQP 6.94E-03 5.05E-01 1.56E-04 3.67E-03 1.95E-01
MNLI 6.94E-03 4.89E-01 1.73E-04 3.57E-03 2.27E-01
QNLI 6.94E-03 4.62E-01 1.53E-04 3.36E-03 2.23E-01
RTE 6.94E-03 5.05E-01 2.15E-04 3.72E-03 2.88E-01
Table 8: Orthogonality analysis using $\text{LLaMA-2}_{\text{7B}}$ on 5 tasks.
Task $\tilde{\text{u}}] E$$\tilde{\text{v}}$] Cov($\tilde{\text{u}}$$\tilde{\text{v}}$) E[$\tilde{\text{u}}$$\tilde{\text{v}}$⟩] Correlation
HellaSwag 9.77E-04 7.56E-01 -1.47E-04 5.92E-04 -1.44E-01
Winogrande 9.77E-04 7.66E-01 -1.85E-04 5.63E-04 -1.66E-01
ARC-e 9.77E-04 7.76E-01 -1.42E-04 6.16E-04 -1.44E-01
ARC-c 9.77E-04 7.71E-01 -1.27E-04 6.26E-04 -1.33E-01
OBQA 9.77E-04 8.01E-01 -1.29E-04 6.53E-04 -1.61E-01
D.2 Seed Sensitivity and Statistical Significance
We evaluate the seed sensitivity of HIS and HIES, with results reported in Table 9 and Table 10 .
Table 9: Accuracy and standard deviation in accuracy across 5 random seeds for $\text{BERT}_{\text{base}}$ on GLUE benchamrk.
Pruning Ratio (%) HIS HIES (ours)
Accuracy (%) Standard Deviation Accuracy (%) Standard Deviation
CoLA 10 74.18 1.60E-02 75.85 1.21E-02
20 74.18 5.44E-04 73.04 1.09E-02
30 72.70 6.91E-03 75.17 2.82E-03
40 70.28 5.52E-03 70.20 3.68E-02
50 60.37 2.18E-02 71.17 5.29E-03
MNLI 10 84.13 0.00E+00 84.06 1.64E-01
20 82.93 9.13E-02 82.87 0.00E+00
30 82.40 0.00E+00 82.29 3.58E-01
40 81.46 2.39E-01 81.39 1.98E-01
50 79.47 4.34E-01 78.87 3.27E-01
MRPC 10 90.29 4.58E-03 90.97 8.04E-04
20 89.69 4.50E-03 89.17 3.55E-03
30 88.79 2.36E-03 89.52 3.69E-03
40 88.02 7.41E-03 87.12 9.06E-03
50 86.85 1.65E-02 87.19 1.66E-02
QNLI 10 90.55 8.69E-02 90.51 1.34E-01
20 90.23 3.48E-01 90.05 2.73E-01
30 88.33 7.47E-01 88.15 7.43E-01
40 84.75 1.45E+00 85.02 8.90E-01
50 80.81 3.51E+00 82.83 1.74E+00
QQP 10 90.68 2.33E-02 91.26 8.00E-03
20 90.56 1.52E-01 90.56 7.31E-02
30 89.84 3.67E-01 89.72 3.45E-01
40 88.07 1.84E-01 87.96 2.54E-01
50 85.33 9.93E-01 85.90 6.97E-01
RTE 10 72.42 2.89E-01 72.13 2.70E-01
20 71.91 5.78E-01 70.76 9.69E-01
30 71.26 1.18E+00 71.34 1.47E+00
40 67.94 2.10E+00 66.79 1.92E+00
50 65.99 1.36E+00 65.42 1.56E+00
SST-2 10 91.86 1.62E-01 92.64 4.59E-02
20 91.77 2.22E-01 92.11 1.69E-01
30 91.19 3.03E-01 92.20 1.92E-01
40 88.67 1.62E+00 90.96 7.16E-01
50 86.31 1.35E+00 90.55 1.24E+00
STS-B 10 87.85 1.88E-03 87.95 3.37E-03
20 87.33 3.52E-03 87.35 2.06E-04
30 86.82 2.15E-03 86.76 3.25E-03
40 85.89 4.49E-03 85.87 4.41E-03
50 83.38 8.95E-03 82.27 1.37E-02
Avg. 10 85.25 7.30E-02 85.67 7.98E-02
20 84.82 1.75E-01 84.49 1.87E-01
30 83.92 3.26E-01 84.39 3.90E-01
40 81.89 7.00E-01 81.91 5.04E-01
50 78.56 9.62E-01 80.53 6.99E-01
Table 10: Accuracy and standard deviation in accuracy across 5 random seeds for $\text{LLaMA-2}_{\text{7B}}$ on HellaSwag, Winogrande, and ARC-c.
Pruning Ratio (%) HIS HIES (ours)
Accuracy (%) Standard Deviation Accuracy (%) Standard Deviation
[…]
D.3 Effect of Calibration Dataset Size
We analyze the effect of calibration dataset size on the performance and stability of HIES. Tables 11 and 12 report accuracy and standard deviation across calibration sizes ranging from 1 to 1024 samples for $\text{BERT}_{\text{base}}$ and $\text{LLaMA-2}_{\text{7B}}$, respectively. Across both model families, HIES exhibits stable performance once a modest number of calibration samples is used. In particular, increasing the calibration set size beyond small values yields diminishing returns in accuracy, while the variance across random seeds remains low. These results indicate that HIES does not require large calibration datasets to obtain reliable head-level statistics, supporting its practical applicability in low-cost calibration settings.
Table 11: Accuracy and standard deviation in accuracy for different calibration dataset sizes (1, 16, 32, 64, 128, 512, 1024) for $\text{BERT}_{\text{base}}$ on GLUE Benchmarks. Note that we use a default size of 32 in the main results.
Pruning Ratio (%) HIS HIES (ours)
Accuracy (%) Standard Deviation Accuracy (%) Standard Deviation
CoLA 10 74.46 1.33E-02 75.71 1.34E-02
30 74.09 8.40E-03 75.10 3.04E-03
50 62.01 3.37E-02 71.04 1.13E-02
MNLI 10 84.13 0.00E+00 84.13 0.00E+00
30 82.40 0.00E+00 82.35 3.48E-01
50 78.93 5.72E-01 78.65 1.98E-01
MRPC 10 90.52 2.82E-03 91.05 2.08E-03
30 89.12 5.97E-03 89.61 2.94E-03
50 86.45 1.55E-02 86.92 1.44E-02
QNLI 10 90.51 9.57E-02 90.54 1.25E-01
30 88.25 5.97E-01 88.01 5.53E-01
50 80.50 2.96E+00 81.97 1.39E+00
QQP 10 90.74 7.25E-02 91.24 4.16E-02
30 90.12 2.31E-01 90.01 2.37E-01
50 84.75 6.91E-01 85.60 4.29E-01
RTE 10 72.36 4.09E-01 72.25 3.25E-01
30 71.17 1.13E+00 70.81 1.13E+00
50 65.45 1.14E+00 65.50 2.13E+00
SST-2 10 91.79 1.30E-01 92.56 1.68E-01
30 90.78 4.53E-01 92.12 2.06E-01
50 85.76 2.30E+00 90.33 1.16E+00
STS-B 10 87.98 7.08E-04 87.96 2.88E-03
30 86.87 2.03E-03 86.75 2.64E-03
50 83.53 6.39E-03 82.80 1.37E-02
Avg. 10 85.31 9.06E-02 85.68 8.48E-02
30 84.10 3.04E-01 84.34 3.10E-01
50 78.42 9.63E-01 80.35 6.68E-01
Table 12: Accuracy and standard deviation in accuracy for different calibration dataset sizes (1, 16, 32, 64, 128, 512, 1024) for $\text{LLaMA-2}_{\text{7B}}$ on HellaSwag, Winogrande, and ARC-c. Note that we use a default size of 32 in the main results.
Pruning Ratio (%) HIS HIES (ours)
Accuracy (%) Standard Deviation Accuracy (%) Standard Deviation
[…]
D.4 Heatmap of Importance Scores and Pruning Results